A projectile can have the same range for two angle of projection . if times of flight in two cases are t1 and t2 then velocity of projection of the projectile is

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
let the velocity is u.
the 2 angle of projection are  and 90-
t1=2ucos/g
t2=2usin/g
t12+t22 = 4u2/g2 (since sin2+cos2 = 1)
=>u=(g/2)√(t1^2+t2^2 )
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)